OpenStudy (anonymous):

Find the common ratio, r, in the following geometric sequence. 27, 9, 3, 1, …

OpenStudy (anonymous):

to get common ratio you can divide number with its pervious one, it will give you common ratio. like 9/27 = 1/3 , 3/9 = 1/3 so , from above we can see that common ration is same so its is a G.P series. so next term can be find just multiply common ratio with last term.